# Iterated sampling importance resampling with adaptive number of proposals

### v1

@@ 1. Introduction
We present an adaptive i-SIR algorithm, which automatically selects the number of proposals $N$, which minimises the cost weighted approximate asymptotic variance [glynn-whitt]. We start by presenting the i-SIR and its continuous-number-of-proposals relaxation (Section 2) and summarise some its theoretical properties (Section 3). In particular, Theorem 4 guarantees that the asymptotic variance of i-SIR is strictly decreasing and convex with respect to the number of proposals. We then discuss its efficiency and present an approximate efficiency in Section 4, which our adaptive algorithm discussed in Section 5 aims to minimise. We prove a strong law of large numbers for our adaptive i-SIR algorithm (Theorem 7), which guarantees consistency of estimates based on its output.
Sections 69 are devoted to technical results required by our theoretical results. Section 6 is about the behaviour of stationary covariances and acceptance rates, which lead to the convexity of the asymptotic variance in Section 7. Section 8 presents upper and lower bounds for i-SIR acceptance rate and asymptotic variance and Section 9 discusess i-SIR in discrete state spaces, and its alternative approximation.
@@ 2. The i-SIR algorithm
The i-SIR algorithm defines a Markov chain Monte Carlo method targetting a probability density $\pi(x)$. It uses samples from a proposal probability density $q(x)$, which is assumed to dominate $\pi$, so that the (unnormalised) importance weights $w(x)=c\pi(x)/q(x)\in[0,\infty)$ are well-defined for all $x$, where $c>0$ is a constant, whose value need not be known; typically it stems from that we can evaluate only the unnormalised target density $\pi_{u}(x)=c\pi(x)$, in which case $c=(\int\pi_{u}(x)\mathrm{d}x)^{-1}$ is the unknown normalising constant.
@@
2:With probability $\beta_{k}=\bar{N}_{k}-\lambda$ set $N_{k}=\bar{N}_{k}-1$, else set $N_{k}=\bar{N}_{k}$. and calculate weights $W_{k}^{i}=w(X_{k}^{i})/\sum_{j=1}^{N_{k}}w(X_{k}^{j})$
@@ 3. Properties of the i-SIR
It is well-known that the i-SIR algorithm $P_{N}$ is reversible with respect to $\pi$ [e.g. tjelmeland, andrieu-lee-vihola], and therefore the generalised i-SIR $P_{\lambda}$, as a convex combination of $P_{N}$ and $P_{N+1}$ is $\pi$-reversible. We assume the following to hold throughout Sections 35.
@@
The i-SIR transition probability $P_{\lambda}$ satisfies the following strong minorization condition:
@@
Fix $\lambda\in[2,\infty)$ and assume that $(X_{k})_{k\geq 0}$ is the stationary i-SIR chain following $P_{\lambda}$, that is, $X_{0}\sim\pi$ and $X_{k}\mid(X_{k-1}=x)\sim P_{\lambda}(x,\;\cdot\;)$. The asymptotic variance of $f\in L^{2}(\pi)$ is:
@@
Let $X_{k}$ be the i-SIR chain with an arbitrary initial state $X_{0}=x$. Then, for any $f\in L^{2}(\pi)$:
@@
Let $f\in L^{2}(\pi)$ such that $\mathrm{var}_{\pi}(f)>0$. Then,
@@
    $\lambda\mapsto\mathrm{var}(P_{\lambda},f)$ is strictly convex and decreasing.
@@ 4. Overall efficiency and approximation based on average acceptance rate
The upper bound of the asymptotic variance (i) might be a possible proxy to optimise. However, while the existence of $\hat{w}$$\infty$ in (A1) might be known, its value is rarely known. It is also unclear how well the upper bound matches with the actual asymptotic variance. (We present some numerical comparisons in Section 10 regarding this.)
@@
Proposition 6 (i) follows from Lemma 37 in Appendix A, (ii) from Lemma 29 and (iii) from Lemma 24. Proposition 6 indicates that $\mathrm{var}(\tilde{P}_{\lambda},f)$ has similar properties as $\mathrm{var}(P_{\lambda},f)$ (cf. Theorem 4). Our method for finding an appropriate ‘number of proposals’ $\lambda$ is based on this expression: we aim for approximately minimising $c(\lambda)\mathrm{var}(\tilde{P}_{\lambda},f)$, which we discuss next.
@@ 5. Adaptation of the number of proposals
$\epsilon(\lambda)=\int\pi(z_{1})\bigg(\prod_{i\geq 2}q(z_{i})\bigg)w(z_{1})\bigg(\frac{\beta(\lambda)}{\sum_{j=1}^{\lfloor\lambda\rfloor}w(z_{j})}+\frac{1-\beta(\lambda)}{\sum_{j=1}^{\lfloor\lambda\rfloor+1}w(z_{j})}\bigg)\mathrm{d}z_{1}\cdots\mathrm{d}z_{n},$
where $\beta(\lambda)=\lfloor\lambda\rfloor+1-\lambda$. Clearly, the function $\epsilon$ is differentiable at non-integer points $\lambda$:
(2) $\epsilon^{\prime}(\lambda)$ $=\int\pi(z_{1})\bigg(\prod_{i\geq 2}q(z_{i})\bigg)w(z_{1})\bigg(\frac{1}{\sum_{j=1}^{\lfloor\lambda\rfloor+1}w(z_{j})}-\frac{1}{\sum_{j=1}^{\lfloor\lambda\rfloor}w(z_{j})}\bigg)\mathrm{d}z_{1}\cdots\mathrm{d}z_{n},$
and even though the derivative is not defined on integer points, the càdlàg function on the right hand side of (2) is. Hereafter we denote by $\epsilon^{\prime}$ the extended derivative function. If we iterated Algorithm 2 (with fixed $\lambda$), we could estimate $\epsilon(\lambda)$ and $\epsilon^{\prime}(\lambda)$ by averaging the following quantities:
@@
The algorithm uses stochastic observations, whose stationary expectations coincide with $G$. It is in instance of stochastic approximation type adaptive MCMC [cf. andrieu-robert, andrieu-moulines, andrieu-thoms] where $\gamma_{k}$ is a constant step size sequence, which we choose as $k^{-\beta}$ where $\beta=0.75$; values $\beta\in(1/2,1)$ are advisable [e.g. andrieu-moulines]. The additional projection function $\mathrm{Proj}\{\xi\}=\min\big\{\max\{0,\xi\},\log(N_{\max}-1)\big\}$ ensures that $2\leq\lambda_{k}\leq N_{\max}<\infty$. The upper bound is added to guarantee the stability of the adaptation, but also simplifies the practical implementation of the method. The initial value of $\xi_{0}$ is arbitrary, and can correspond for instance to $\lambda_{0}=N_{\max}/2$.
@@
Suppose that $\gamma_{k}=k^{-\beta}$ where $\beta>0$ and the cost function $c:[2,N_{\max}]\to(0,\infty)$ is continuous and differentiable. Then, the variables $X_{1},X_{2},\ldots$ generated by iteration of Algorithm 3 satisfy the following strong law of large numbers for any bounded (measurable) test function $f$:
@@
Let us denote
@@
By definition $c^{\prime}$ is bounded. Additionally $|1-\hat{\epsilon}_{k}(\lambda_{k-1})^{2}|\leq 1$ holds always. Now $\lambda_{k-1}=e^{\xi_{k-1}}+1$, let us denote $N(x)=\lfloor e^{x}+1\rfloor$, then
$\mathbb{E}[|\hat{\epsilon}_{k}^{\prime}(\lambda_{k-1})|\mid X_{k-1},\xi_{k-1}]$ $=\mathbb{E}\bigg[\bigg|\frac{w(Y_{k}^{1})}{\sum_{j=1}^{N(\xi_{k-1})}w(Y_{k}^{j})}-\frac{w(Y_{k}^{1})}{\sum_{j=1}^{N(\xi_{k-1})+1}w(Y_{k}^{j})}\bigg|\;\bigg|\;X_{k-1},\xi_{k-1}\bigg]$
$=\mathbb{E}\bigg[\frac{w(Y_{k}^{1})w(Y_{k}^{N(\xi_{k-1})+1})}{\big(\sum_{j=1}^{N(\xi_{k-1})}w(Y_{k}^{j})\big)\big(\sum_{j=1}^{N(\xi_{k-1})+1}w(Y_{k}^{j})\big)}\;\bigg|\;X_{k-1},\xi_{k-1}\bigg]$
@@
$\mathbb{E}[|c(\lambda_{k-1})||\hat{\epsilon}_{k}^{\prime}(\lambda_{k-1})|\mid X_{k-1},\xi_{k-1}]\leq\frac{|c(e^{\xi_{k-1}}+1)|}{\lfloor e^{\xi_{k-1}}+1\rfloor}\leq\frac{|c(e^{\xi_{k-1}}+1)|}{e^{\xi_{k-1}}},$
which is bounded and thus
$\mathbb{E}[|H_{k}|]\leq\mathbb{E}[|c^{\prime}(\lambda_{k-1})|(1-\hat{\epsilon}_{k}(\lambda_{k-1})^{2}]+2\mathbb{E}[\mathbb{E}[|c(\lambda_{k-1})\hat{\epsilon}^{\prime}_{k}(\lambda_{k-1})|\mid X_{k-1},\xi_{k-1}]]\leq M,$
for some $M>0$. By definition
$\mathrm{Proj}(\xi_{k-1}+\gamma_{k}H_{k})=\begin{cases}\xi_{k-1}+\gamma_{k}H_{k},&\xi_{k-1}+\gamma_{k}H_{k}\in[0,\log(N_{\max}-1)],\\ 0,&\xi_{k-1}+\gamma_{k}H_{k}<0\\ \log(N_{\max})-1,&\xi_{k-1}+\gamma_{k}H_{k}>\log(N_{\max}-1).\end{cases}$
Let us choose $\tilde{H}_{k}=\frac{\mathrm{Proj}(\xi_{k-1}+\gamma_{k}H_{k})-\xi_{k-1}}{\gamma_{k}}$, then we obtain
$\xi_{k}=\mathrm{Proj}(\xi_{k-1}+\gamma_{k}H_{k})=\xi_{k-1}+\gamma_{k}\tilde{H}_{k}$
and
$|\lambda_{k}-\lambda_{k-1}|=|\exp(\xi_{k})-\exp(\xi_{k-1})|\leq L|\xi_{k}-\xi_{k-1}|\leq L\gamma_{k}|\tilde{H}_{k}|,$
where $L$ is Lipschitz constant for $g:[0,\log(N_{\max}-1)]\to\mathbb{R}$, $g(x)=\exp(x)$. Clearly
$\gamma_{k}|\tilde{H}_{k}|\leq\gamma_{k}|H_{k}|$
and thus
$\sum_{k=1}^{\infty}\frac{\gamma_{k}}{k}\mathbb{E}\big[|\tilde{H}_{k}|\big]\leq\sum_{k=1}^{\infty}k^{-(\beta+1)}\mathbb{E}\big[|H_{k}|\big]\leq M\sum_{k=1}^{\infty}k^{-(\beta+1)}<\infty.$
By combining this, Lemma 47 and 49 in Appendix E with [laitinen-vihola-adaptive, Lemma 16 $(ii)$, Theorem 9 $(ii)$] we obtain strong law of large numbers. ∎
@@
For a cost function $c(\lambda)=a+b\lambda$ where $a\geq 0$ and $b>0$ and a test function $L^{2}(\pi)$, $\mathrm{var}_{\pi}(f)>0$, the minimiser $\lambda^{*}$ of $c(\lambda)\mathrm{var}(P_{\lambda},f)$ satisfies $\lambda^{*}\leq 4\sqrt{\hat{w}(a/b+1)}+4\hat{w}+1$ by Proposition 50 in Appendix F.

### v2

@@ 1. Introduction
We present an adaptive i-SIR algorithm, which automatically selects the number of proposals $N$, which minimises the cost weighted approximate asymptotic variance [glynn-whitt]. We start by presenting the i-SIR and its continuous-number-of-proposals relaxation (Section 2) and summarise some of its theoretical properties (Section 3). In particular, Theorem 4 guarantees that the asymptotic variance of i-SIR is strictly decreasing and convex with respect to the number of proposals. We then discuss its efficiency and present an approximate efficiency in Section 4, which our adaptive algorithm discussed in Section 5 aims to minimise. We prove a strong law of large numbers for our adaptive i-SIR algorithm (Theorem 7), which guarantees consistency of estimates based on its output.
Sections 69 are devoted to technical results required by our theoretical results. Section 6 is about the behaviour of stationary covariances and acceptance rates, which lead to the convexity of the asymptotic variance in Section 7. Section 8 presents upper and lower bounds for i-SIR acceptance rate and asymptotic variance, and Section 9 discusses i-SIR in discrete state spaces, and its alternative approximation.
@@ 2. The i-SIR algorithm
The i-SIR algorithm defines a Markov chain Monte Carlo method targeting a probability density $\pi(x)$. It uses samples from a proposal probability density $q(x)$, which is assumed to dominate $\pi(x)$, that is, $q(x)=0$ implies $\pi(x)=0$. Then, the (unnormalised) importance weights $w(x)=c\pi(x)/q(x)\in[0,\infty)$ are well-defined for all $x$, where we use the convention $0/0=0$. The value of the constant $c>0$ need not be known; typically it stems from that we can evaluate only the unnormalised target density $\pi_{u}(x)=c\pi(x)$, in which case $c=(\int\pi_{u}(x)\mathrm{d}x)^{-1}$ is the unknown normalising constant.
@@
2:With probability $\beta_{k}=\bar{N}_{k}-\lambda$ set $N_{k}=\bar{N}_{k}-1$, else set $N_{k}=\bar{N}_{k}$, and calculate weights $W_{k}^{i}=w(X_{k}^{i})/\sum_{j=1}^{N_{k}}w(X_{k}^{j})$
@@ 3. Properties of the i-SIR
It is well-known that the i-SIR algorithm $P_{N}$ is reversible with respect to $\pi$ [e.g. tjelmeland, andrieu-lee-vihola], and therefore the generalised i-SIR $P_{\lambda}$, as a convex combination of $P_{N}$ and $P_{N+1}$ is $\pi$-reversible. We assume the following to hold throughout Sections 35:
@@
The i-SIR transition probability $P_{\lambda}$ satisfies the following strong minorisation condition:
@@
Fix $\lambda\in(1,\infty)$ and assume that $(X_{k})_{k\geq 0}$ is the stationary i-SIR chain following $P_{\lambda}$, that is, $X_{0}\sim\pi$ and $X_{k}\mid(X_{k-1}=x)\sim P_{\lambda}(x,\;\cdot\;)$. The asymptotic variance of $f\in L^{2}(\pi)$ is:
@@
For $\lambda\in(1,\infty)$, let $X_{k}$ be the i-SIR chain with an arbitrary initial state $X_{0}=x$. Then, for any $f\in L^{2}(\pi)$:
@@
Let $\lambda\in(1,\infty)$ and $f\in L^{2}(\pi)$ such that $\mathrm{var}_{\pi}(f)>0$. Then,
@@
    $\lambda\mapsto\mathrm{var}(P_{\lambda},f)$ is strictly convex and decreasing on $(1,\infty)$.
@@ 4. Overall efficiency and approximation based on average acceptance rate
The upper bound of the asymptotic variance (i) might be a possible proxy to optimise. However, while the existence of $\hat{w}$$\infty$ in (A1) might be known, its value is rarely known. It is also unclear how well the upper bound matches the actual asymptotic variance. (We present some numerical comparisons in Section 10 regarding this.)
@@
Proposition 6 (i) follows from Lemma 37 in Appendix B, (ii) from Lemma 29 and (iii) from Lemma 24. Proposition 6 indicates that $\mathrm{var}(\tilde{P}_{\lambda},f)$ has similar properties as $\mathrm{var}(P_{\lambda},f)$ (cf. Theorem 4). Our method for finding an appropriate ‘number of proposals’ $\lambda$ is based on this expression: we aim for approximately minimising $c(\lambda)\mathrm{var}(\tilde{P}_{\lambda},f)$, which we discuss next.
@@ 5. Adaptation of the number of proposals
$\epsilon(\lambda)=\int\pi(z_{1})\bigg(\prod_{i=2}^{N(\lambda)}q(z_{i})\bigg)w(z_{1})\bigg(\frac{\beta(\lambda)}{\sum_{j=1}^{N(\lambda)-1}w(z_{j})}+\frac{1-\beta(\lambda)}{\sum_{j=1}^{N(\lambda)}w(z_{j})}\bigg)\mathrm{d}z_{1}\cdots\mathrm{d}z_{N(\lambda)},$
where $\beta(\lambda)=\lfloor\lambda\rfloor+1-\lambda$ and $N(\lambda)=\lfloor\lambda\rfloor+1$ Clearly, the function $\epsilon$ is differentiable at non-integer points $\lambda$:
(2) $\epsilon^{\prime}(\lambda)$ $=\int\pi(z_{1})\bigg(\prod_{i=2}^{N(\lambda)}q(z_{i})\bigg)w(z_{1})\bigg(\frac{1}{\sum_{j=1}^{N(\lambda)}w(z_{j})}-\frac{1}{\sum_{j=1}^{N(\lambda)-1}w(z_{j})}\bigg)\mathrm{d}z_{1}\cdots\mathrm{d}z_{N(\lambda)},$
and we extend $\epsilon^{\prime}$ to all $\lambda\geq 1$ by taking the cádlág extension of the right-hand side of (2). If we iterated Algorithm 2 (with fixed $\lambda$), we could estimate $\epsilon(\lambda)$ and $\epsilon^{\prime}(\lambda)$ by averaging the following quantities:
@@
The algorithm uses stochastic observations, whose stationary expectations coincide with $G$. It is in instance of stochastic approximation type adaptive MCMC [cf. andrieu-robert, andrieu-moulines, andrieu-thoms] where $\gamma_{k}$ is a constant step size sequence, which we choose as $k^{-\beta}$ where $\beta=0.75$; values $\beta\in(1/2,1)$ are advisable [e.g. andrieu-moulines]. The additional projection function $\mathrm{Proj}\{\xi\}=\min\big\{\max\{0,\xi\},\log(N_{\max}-1)\big\}$ ensures that $2\leq\lambda_{k}\leq N_{\max}$ when $N_{\max}$ is finite, and $2\leq\lambda_{k}<\infty$ when $N_{\max}=\infty$. A finite upper bound is not necessary for the validity, but it can simplify the practical implementation of the method. The initial value of $\xi_{0}$ is arbitrary, and can correspond for instance to $\lambda_{0}=N_{\max}/2$.
@@
Suppose that $\gamma_{k}=k^{-\beta}$ where $\beta>0$. Suppose the cost function $c$ is continuous and differentiable, and in the case $N_{\max}=\infty$, additionally $c(\lambda)=O(\lambda)$ and $c^{\prime}$ is bounded. Then, the variables $X_{1},X_{2},\ldots$ generated by iteration of Algorithm 3 satisfy the following strong law of large numbers for any bounded (measurable) test function $f$:
@@
When $N_{\max}$ is finite, we set $A=[0,\log(N_{\max}-1)]$, and when $N_{\max}=\infty$, we set $A=[0,\infty)$. Let us denote
@@
By definition
$\mathrm{Proj}(\xi_{k-1}+\gamma_{k}H_{k})=\begin{cases}\xi_{k-1}+\gamma_{k}H_{k},&\xi_{k-1}+\gamma_{k}H_{k}\in A,\\ 0,&\xi_{k-1}+\gamma_{k}H_{k}<0\\ \log(N_{\max}-1),&\xi_{k-1}+\gamma_{k}H_{k}>\log(N_{\max}-1).\end{cases}$
Let $\tilde{H}_{k}=\frac{\mathrm{Proj}(\xi_{k-1}+\gamma_{k}H_{k})-\xi_{k-1}}{\gamma_{k}}$, then we obtain
$\xi_{k}=\mathrm{Proj}(\xi_{k-1}+\gamma_{k}H_{k})=\xi_{k-1}+\gamma_{k}\tilde{H}_{k},$
$|\xi_{k}-\xi_{k-1}|=\gamma_{k}|\tilde{H}_{k}|$ and clearly $\gamma_{k}|\tilde{H}_{k}|\leq\gamma_{k}|H_{k}|$. We have $|1-\hat{\epsilon}_{k}(\lambda_{k-1})^{2}|\leq 1$. Now $\lambda_{k-1}=e^{\xi_{k-1}}+1$, let us denote $N(x)=\lfloor e^{x}+1\rfloor$, then
$\mathbb{E}\big[|\hat{\epsilon}_{k}^{\prime}(\lambda_{k-1})|\;\big|\;X_{k-1},\xi_{k-1}\big]$ $=\mathbb{E}\Bigg[\bigg|\frac{w(Y_{k}^{1})}{\sum_{j=1}^{N(\xi_{k-1})+1}w(Y_{k}^{j})}-\frac{w(Y_{k}^{1})}{\sum_{j=1}^{N(\xi_{k-1})}w(Y_{k}^{j})}\bigg|\;\Bigg|\;X_{k-1},\xi_{k-1}\Bigg]$
$=\mathbb{E}\bigg[\frac{w(Y_{k}^{1})w(Y_{k}^{N(\xi_{k-1})+1})}{\big(\sum_{i=1}^{N(\xi_{k-1})}w(Y_{k}^{i})\big)\big(\sum_{j=1}^{N(\xi_{k-1})+1}w(Y_{k}^{j})\big)}\;\bigg|\;X_{k-1},\xi_{k-1}\bigg]$
@@
$\mathbb{E}\big[|c(\lambda_{k-1})\hat{\epsilon}_{k}^{\prime}(\lambda_{k-1})|\;\big|\;X_{k-1},\xi_{k-1}\big]\leq\frac{|c(e^{\xi_{k-1}}+1)|}{\lfloor e^{\xi_{k-1}}+1\rfloor}\leq\frac{|c(e^{\xi_{k-1}}+1)|}{e^{\xi_{k-1}}}.$
When $N_{\max}$ is finite, functions $x\mapsto\frac{|c(e^{x}+1)|}{e^{x}}$ and $x\mapsto c^{\prime}(e^{x})$ are bounded on $A$. When $N_{\max}=\infty$ and the additional assumptions hold, then $x\mapsto\frac{|c(e^{x}+1)|}{e^{x}}$ is continuous and $\frac{|c(e^{x}+1)|}{e^{x}}=O(1)$, hence bounded on $A$. Thus, in both cases
$\mathbb{E}\big[|H_{k}|\big]\leq\mathbb{E}\big[|c^{\prime}(\lambda_{k-1})|(1-\hat{\epsilon}_{k}(\lambda_{k-1})^{2}\big]+2\mathbb{E}\big[\mathbb{E}\big[|c(\lambda_{k-1})\hat{\epsilon}^{\prime}_{k}(\lambda_{k-1})|\;\big|\;X_{k-1},\xi_{k-1}\big]\big]\leq M,$
for some $M>0$, hence
$\sum_{k=1}^{\infty}\frac{\gamma_{k}}{k}\mathbb{E}\big[|\tilde{H}_{k}|\big]\leq\sum_{k=1}^{\infty}\frac{\gamma_{k}}{k}\mathbb{E}\big[|H_{k}|\big]=\sum_{k=1}^{\infty}k^{-(\beta+1)}\mathbb{E}\big[|H_{k}|\big]\leq M\sum_{k=1}^{\infty}k^{-(\beta+1)}<\infty.$
Let $\xi,\xi^{\prime}\in A$, then by Lemma 49
$d_{\mathrm{tv}}\big(P_{e^{\xi^{\prime}}+1}(x,\;\cdot\;),P_{e^{\xi}+1}(x,\;\cdot\;)\big)$ $\leq\frac{6|e^{\xi^{\prime}}-e^{\xi}|}{\max\{e^{\xi^{\prime}}+1,e^{\xi}+1\}}\leq 6|1-e^{-|\xi^{\prime}-\xi|}|\leq 6|\xi^{\prime}-\xi|,$
where the last inequality follows since $x\mapsto x-(1-e^{-x})$ is non-negative and increasing on $0,\infty)$. Therefore, by combining all the previous and Lemma 50 in Appendix E with [laitinen-vihola-adaptive, Lemma 16 $(ii)$, Theorem 9 $(ii)$], we obtain strong law of large numbers. ∎
@@
For a cost function $c(\lambda)=a+b\lambda$ where $a\geq 0$ and $b>0$ and a test function $f\in L^{2}(\pi)$, $\mathrm{var}_{\pi}(f)>0$, the minimiser $\lambda^{*}$ of $c(\lambda)\mathrm{var}(P_{\lambda},f)$ satisfies $2\leq\lambda^{*}\leq 4\sqrt{\hat{w}(a/b+1)}+4\hat{w}+1$ by Proposition 51 and 54 in Appendix F.